OpenStudy (anonymous):

also looking for the X in this sequence: 7-5-4-8-6-3-15-11-X there is a logic to it but I can't find it and also this sequence: 992-101-496-55-X-32-124-16

OpenStudy (anonymous):

Should be 7 for the first sequence.

OpenStudy (anonymous):

7+8=15 5+6=11 4+3=7

OpenStudy (anonymous):

Is the second number in the second sequence supposed to be 110?

OpenStudy (anonymous):

no it's 101...

OpenStudy (anonymous):

If it is, X should be 248.

OpenStudy (anonymous):

The secondary sequence doesn't make sense. The paired sets 992/2 =496 496/2=248 248/2=124

OpenStudy (anonymous):

101-55-32-16 I can't do much with those. 110 would be 55*2 and 32 is 16*2.

OpenStudy (anonymous):

that was my first idea as well but in all other sequences all numbers are used... with this logic it could be an error in the question...

OpenStudy (anonymous):

I'd lean towards a typo, but there is the gap from 55 to 32 as well.

OpenStudy (anonymous):

Maybe someone else can find something. Sorry :)

OpenStudy (anonymous):

thanks for trying!